using Ax register convert this high-level language expression to Assembly language code

- C=C+4

- C=C-1
Answer #1

EXPRESSION: C= C+4

Assembly Code:

  • MOV Ax, C
  • ADD Ax, #4

EXPLANATION:

We are given C=C+4

So to write this high level language code into assmebly code:

Let's first mov the content of variable 'C' to Register 'Ax' by MOV Ax, C

Where Ax is destination and C is source, so content is moved from source to destination register.

And then, add 4 to the content in Register Ax

For that we have ADD instruction.

-----------------------------------------------------------------------------

EXPRESSION: C=C-1

Assembly Code:

  • MOV Ax, C
  • DEC Ax

EXPLANATION:

Here we have expression as: C=C-1

So first mov the content of C to Register 'Ax'

With the help of instruction MOV Ax, C

Now subtract 1 from the content of Ax

With the help of help of DEC instruction, which means to decrement the content of register by 1.

Hence, DEC AX